Solusi dari persamaan sin xdx + ydy = 0; y(0)=-2
1
Expert's answer
2020-12-04T09:58:04-0500

Let us solve the equation "\sin xdx + ydy = 0; y(0)=-2":


"\sin xdx =- ydy"


"\int \sin xdx = -\int ydy"


"-\cos x=-\frac{y^2}{2}+C"


Since "y(0)=-2", we have the equation "-\cos 0=-\frac{(-2)^2}{2}+C" which is equivalent to "-1=-2+C". So, "C=1."


Therefore,


"-\cos x=-\frac{y^2}{2}+1" or


"y^2=2\cos x+2"


Since "y(0)=-2", the solution is


"y=-\sqrt{2\cos x+2}"